OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle (x + 6)2 + (y + 4)2 = 25 at the point (-9, -8). y=4/3x−593 y=−34x+593 y=4/3x+594 y=−3/4x−594

OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle (x + 6)2 + (y + 4)2 = 25 at the point (-9, -8). y=4/3x−59/3 y=−3/4x+59/3 y=4/3x+59/4 y=−3/4x−59/4

ganeshie8 (ganeshie8):

You need a `point` and `slope` to write the equation of line

ganeshie8 (ganeshie8):

you already have a `point` (-9, -8)

ganeshie8 (ganeshie8):

any ideas how to find the slope of the required tangent ?

OpenStudy (anonymous):

the 593 are suposeed to be 59/3 and 59/4

OpenStudy (anonymous):

no i domnt im confused

ganeshie8 (ganeshie8):

see if you can use the fact that \(\large \text{tangent} \perp \text{radius}\) \

ganeshie8 (ganeshie8):

start by finding the slope of that radius segment

ganeshie8 (ganeshie8):

it ends in : (-6, -4) and (-9, -8) slope = ?

OpenStudy (anonymous):

i got y= 3/4x - 59/4

OpenStudy (anonymous):

D.

ganeshie8 (ganeshie8):

doesnt look correct

OpenStudy (anonymous):

either that or B

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

did you find the slope of radius segment yet ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

interesting, how did you figure out its either D or B ?

OpenStudy (anonymous):

im guessing good guesser

ganeshie8 (ganeshie8):

guessing is okay if it works

OpenStudy (anonymous):

yes i got it right my answer is indeed correct sir

ganeshie8 (ganeshie8):

okay good for you